OpenStudy (anonymous):

A card is randomly selected from a standard deck. What is the probability that a club or a red queen is selected? a 1/26 b 14/52 c 15/52 d 17/52

Directrix (directrix):

There are 13 clubs in the deck. Of the queens in the deck, 2 are red. Prob(Club or red queen) = 15/52

OpenStudy (anonymous):

probability of a club, A, K, Q, J, 2, 3, 4,....10, would be 13/52 probability of a red queen, diamond or heart, would be 2/52 therefore 13/52 + 2/52 = 15/52 on "or" probability problems, one adds the two probabilities together

OpenStudy (anonymous):

thanks! both of u!!